Suppose we would add a horizontal analyzer D (pointing along the x-axis) after Analyzer C to the experiment shown below. What would be the probability for the atom to exit from either channel? Briefly explain.

Suppose we would add a horizontal analyzer D (poin
Answer #1

ignore gnore

(a) Ignore the atoms exiting from the minus channel of Analyzer 1 and feed the atoms exiting from the plus channel into Analyzer 2.

(b) Ignore the atoms exiting from the minus channel of Analyzer 2 and feed the atoms exiting from the plus channel into Analyzer 3.

When these atoms enter into Analyzer 3, then we get -

Half the atoms leave from the plus-channel of Analyzer 3 and half leave from the minus-channel.

The probability for an atom leaving the plus-channel of Analyzer 2 to exit from the plus-channel of Analyzer 3 which will be given as :

using a formula, we have

P [ x ] = cos2 (\theta/2)

where, \theta = 90 degree

P [ x ] = cos2 (45)0

P [ x ] = 0.5   or 50%